OpenStudy (anonymous):

Write the equation of the line that is perpndicular to the line y = -2x -4 and passes through the point (1, -3)

OpenStudy (luigi0210):

When finding a perpendicular equation you use the negative reciprocal of the slope

OpenStudy (luigi0210):

So do you know the slope?

OpenStudy (anonymous):

-2x

OpenStudy (luigi0210):

Well that's the original.. but the perpendicular slope is 1/2

OpenStudy (anonymous):

oh

OpenStudy (luigi0210):

so use y-y1=m(x-x1)

OpenStudy (anonymous):

y-(-3) = 1/2(x-1)

OpenStudy (luigi0210):

Correct! Now simplify it and write the final equation

OpenStudy (anonymous):

0= 1/2x -3 1/2